# A look at the operator product expansion in critical dynamics

### v1

@@ Appendix B Fokker-Planck Hamiltonian, ordering, and mid-point prescription
$\partial_{t}p\left(x,t\right)$ $=$ $-\partial_{x}\left(A\left(x\right)p\left(x,t\right)\right)+D\partial_{x}^{2}p\left(x,t\right)\,.$
@@
$\partial_{t}p\left(x,t\right)$ $=$ $\left[-i\hat{p}A\left(\hat{x}\right)-D\hat{p}^{2}\right]p\left(x,t\right)$
$=$ $-\left[i\hat{p}A\left(\hat{x}\right)+D\hat{p}^{2}\right]p\left(x,t\right)\,.$
@@
$\left[\hat{p}A\left(\hat{x}\right)\right]_{{\rm Weyl}}$$\equiv$ $\frac{\hat{p}A\left(\hat{x}\right)+A\left(\hat{x}\right)\hat{p}}{2}\,.$
@@
$\partial_{t}p\left(x,t\right)$ $=$ $-\left[i\left[\hat{p}A\left(\hat{x}\right)\right]_{{\rm Weyl}}+\frac{1}{2}A^{\prime}\left(\hat{x}\right)+D\hat{p}^{2}\right]p\left(x,t\right)\,.$
@@
$f\left(x\left(t_{f}\right)\right)-f\left(x\left(t_{0}\right)\right)\,=\,\sum_{j}\left[f\left(x_{j+1}\right)-f\left(x_{j}\right)\right]$ $=$
$\sum_{j}f^{\prime}\left(x_{j}+\lambda\delta\right)\delta+\sum_{j}\frac{1}{2}\left(1-2\lambda\right)f^{\prime\prime}\left(x_{j}\right)\delta^{2}+O\left(\delta^{3}\right)\,.$

### v2

@@ Appendix C Fokker-Planck Hamiltonian, ordering, and mid-point prescription
$\partial_{t}p\left(x,t\right)$ $=$ $-\partial_{x}\left(A\left(x\right)p\left(x,t\right)\right)+D\partial_{x}^{2}p\left(x,t\right)\,.$ (103)
@@
$\partial_{t}p\left(x,t\right)$ $=$ $\left[-i\hat{p}A\left(\hat{x}\right)-D\hat{p}^{2}\right]p\left(x,t\right)$ (104)
$=$ $-\left[i\hat{p}A\left(\hat{x}\right)+D\hat{p}^{2}\right]p\left(x,t\right)\,.$ (105)
@@
$\left[\hat{p}A\left(\hat{x}\right)\right]_{{\rm Weyl}}$$\equiv$ $\frac{\hat{p}A\left(\hat{x}\right)+A\left(\hat{x}\right)\hat{p}}{2}\,.$ (106)
@@
$\partial_{t}p\left(x,t\right)$ $=$ $-\left[i\left[\hat{p}A\left(\hat{x}\right)\right]_{{\rm Weyl}}+\frac{1}{2}A^{\prime}\left(\hat{x}\right)+D\hat{p}^{2}\right]p\left(x,t\right)\,.$ (107)
@@
$f\left(x\left(t_{f}\right)\right)-f\left(x\left(t_{0}\right)\right)\,=\,\sum_{j}\left[f\left(x_{j+1}\right)-f\left(x_{j}\right)\right]$ $=$ (109)
$\sum_{j}f^{\prime}\left(x_{j}+\lambda\delta\right)\delta+\sum_{j}\frac{1}{2}\left(1-2\lambda\right)f^{\prime\prime}\left(x_{j}\right)\delta^{2}+O\left(\delta^{3}\right)\,.$ (110)
@@
Appendix D Action in the large $N$ limit
In this section we rewrite the functional integral associated with the dynamical $O(N)$ model in a way that allows to take the large $N$ limit straightforwardly, see Moshe and Zinn-Justin (2003). The partition function takes the form
$Z$ $=$ $\int{\cal D}\phi{\cal D}\bar{\phi}\exp\Biggr{[}-\int_{tx}\bar{\phi}^{a}\left(\partial_{t}\phi^{a}+D_{0}\left(-\partial^{2}+r_{0}+\frac{u_{0}}{3!N}\phi^{m}\phi^{m}\right)\phi^{a}\right)+\int_{tx}D_{0}\bar{\phi}^{a}\bar{\phi}^{a}\Biggr{]}$ (111)
×$\times\exp\Biggr{[}\frac{1}{2}D\frac{\delta^{2}H}{\delta\phi^{a}\delta\phi^{a}}\Biggr{]}\,$
$\exp\Biggr{[}\frac{1}{2}D\frac{\delta^{2}H}{\delta\phi^{a}\delta\phi^{a}}\Biggr{]}$ $=$ $\exp\Biggr{[}\frac{1}{2}D\int_{t,xy}\left(N\left(-\partial_{x}^{2}+r\right)+\frac{u_{0}}{3!}\left(\phi^{2}+\frac{2}{N}\phi^{2}\right)\right)\delta\left(x-y\right)\Biggr{]}\,.$ (112)
$Z$ $=$ $\int{\cal D}\phi{\cal D}\bar{\phi}\exp\Biggr{[}-\int_{tx}\bar{\phi}^{a}\left(\partial_{t}\phi^{a}+D_{0}\left(-\partial^{2}+r_{0}+\frac{u_{0}}{3!N}\phi^{m}\phi^{m}\right)\phi^{a}\right)+\int_{tx}D_{0}\bar{\phi}^{a}\bar{\phi}^{a}\Biggr{]}$ (113)
$S$ $=$ $\int_{tx}\bar{\sigma}\left(\partial_{t}\sigma+D_{0}\left(-\partial^{2}+r_{0}+\frac{u_{0}}{3!N}\rho\right)\sigma\right)$ (114)
$S$ $=$ $N\Biggr{\{}\int_{tx}\bar{\sigma}\left(\partial_{t}\sigma+D_{0}\left(-\partial^{2}+r_{0}+\frac{u_{0}}{3!}\rho\right)\sigma\right)$ (115)